# Provable Mixed-Noise Learning with Flow-Matching

### v1

@@ 3.3 Gradient Descent Based Framework
Under regularity assumptions of (A1) we can switch the following differentiation and integration, and applying subsequently (A2), we get
$\nabla_{1}^{2}\mathcal{Q}(\theta^{\ast},\theta^{\ast})=\mathbb{E}_{y\sim p_{\theta^{*}}}\mathbb{E}_{x\sim P_{X|Y_{\theta^{*}}=y}}\left[\nabla_{\theta}^{2}\log p_{X,Y_{\theta^{\ast}}}(x,y)\right]\geq-\lambda_{2}\mathrm{Id}.$ (7)
@@ 3.4 Step Size Choice Revisited
$\theta^{(k+1)}=\theta^{(k)}+\tau\nabla q(\theta^{(k)}),$
@@
$0<\sqrt{1-\tau c}+\tau\gamma<1,\quad c\coloneqq\frac{2\mu\lambda}{\lambda+\mu}.$ (8)
@@
For any $\mu,\lambda>0$ and $\gamma$ such that $0\leq\gamma<c=\frac{2\mu\lambda}{\mu+\lambda}$, there exists a step size $\tau\in(0,c^{-1})$ such the gradient ascent iterates in the EM-Algorithm convergence to $\theta^{\ast}$.
@@
If we additionally assume $\mu\geq\lambda$, then $c\geq\lambda,$ with equality if and only if $\mu=\lambda$. In particular in the case $\mu>\lambda$, we can relax the constraints on $\gamma$ from [2], i.e. $0\leq\gamma<\lambda$, for convergence to the range $0\leq\gamma<c$ with $c$ from (8). The resulting improvement is illustrated in Figure˜1.
@@ 4 Verification of Assumptions for Mixed Noise Model
  • Regularity and growth of the forward model: Each component $f_{i}:\mathbb{R}^{d_{x}}\to\mathbb{R}$ is continuously differentiable and satisfies the bounded gradient conditions:
    $\|\nabla f_{i}(x)\|\leq C_{f^{\prime}}\quad\text{for all }x\in\mathbb{R}^{d_{x}}.$
  • Moment bounds: Existence of first and second moment in $Y$-space under the joint distribution.
  • Prior regularity: The prior density $p_{X}$ is Lipschitz continuous and has finite moments up to order 4:
@@
the log likelihood is given (up to an additive constant) by
@@
$=-\frac{1}{2}\left(\sum_{i=1}^{n}\log\sigma_{i}(x)+\frac{r_{i}(x,y)^{2}}{\sigma_{i}(x)}\right).$
@@
$\log p_{Y_{\theta}|X=x}(y)=-\frac{1}{2}\sum_{i=1}^{n}\left(\log\sigma_{i}(x)+\frac{r_{i}(x,y)^{2}}{\sigma_{i}(x)}\right),$
@@
We assume hat the required moments exist under the distribution $p_{X|Y_{\theta}=y}\,p_{\theta^{*}}$.
Since the log-likelihood is smooth in $\theta\in\Theta$, and the derivatives involve rational expressions in $f_{i}(x)$, $r_{i}(x,y)$, and $\sigma_{i}(x)$, all of which are smooth functions of $x$ and $y$, the gradient and Hessian are continuous. Furthermore, under moment bounds on $y$, the gradient and Hessian norms are integrable, since all involved components are uniformly bounded in $x$ as rational polynomials in $f_{i}(x)$ with nominator of degree larger or equal to the nominators. Therefore, the conditions of Assumption (A1)—namely, differentiability of $\theta\mapsto\log p_{X,Y_{\theta}}(x,y)$ and integrable bounds on its gradient and Hessian—are satisfied.
@@
$p_{Y|X=x}(y)=\mathcal{N}(\bm{f}(x),\bm{\Sigma}(x)),\quad\bm{f}(x)=(f_{i}(x))_{i=1}^{n},\quad\bm{\Sigma}(x)=\operatorname{diag}(\sigma_{1}(x),\dots,\sigma_{n}(x)).$
@@
$\mathcal{I}(\theta)=\frac{1}{2}\sum_{i=1}^{n}\int_{\mathbb{R}^{d_{x}}}\frac{1}{\sigma_{i}(x)^{2}}\begin{pmatrix}1&f_{i}(x)^{2}\\ f_{i}(x)^{2}&f_{i}(x)^{4}\end{pmatrix}p_{X}(x)\,\mathrm{d}x.$
@@
Continuity:
Since the entries of the Fisher matrix depend continuously on $\theta$ via $\sigma_{i}(x)=a^{2}+b^{2}f_{i}(x)^{2}$, and both $a^{2}$ and $b^{2}$ are bounded away from zero on the compact domain $\Theta$, the mapping $\theta\mapsto\nabla^{2}_{\theta}\log p_{X,Y_{\theta}}(x,y)$ is continuous on $\Theta$.
@@
To verify Assumption (A3), we follow the stability result in [25, Theorem 14], which gives Lipschitz continuity of the posterior $\theta\mapsto P_{X|Y_{\theta}=y}$ in the Wasserstein-1 distance.
@@
Since $a^{2}>0$, there exists $C_{a^{2}}=C_{a^{2}}(y_{i},a^{2},b^{2})>0$ such that for all $z\in\mathbb{R}$,
$0\leq\left|\frac{1}{(a^{2}+b^{2}z^{2})}-\frac{(y_{i}-z)^{2}}{(a^{2}+b^{2}z^{2})^{2}}\right|<C_{a^{2}}(y_{i},a^{2},b^{2})$
@@
$0\leq\left|z^{2}\left(\frac{1}{(a^{2}+b^{2}z^{2})}-\frac{(y_{i}-z)^{2}}{(a^{2}+b^{2}z^{2})^{2}}\right)\right|<C_{b^{2}}(y_{i},a^{2},b^{2}),$
@@
$\mathrm{maxgrad}\ l(x,y)\leq\sup_{\theta}\frac{1}{2}\sum\limits_{i=1}^{n}\max\{C_{a^{2}}(y_{i},a^{2},b^{2}),C_{b^{2}}(y_{i},a^{2},b^{2})\}=:C_{\mathrm{grad}}(y).$
@@
for $p=1,2$, and some constant $L(y)$ depending polynomial of degree $2$ on $\|y\|$.
Moreover, by [25, Proposition 6], the normalization constants $Z_{1},Z_{2}$ of the posterior densities are bounded from below
$\min(Z_{1},Z_{2})\geq C\exp\left(-\left\|\log p_{Y_{\theta_{1}}|X=x}(y)-\log p_{Y_{\theta_{2}}|X=x}(y)\right\|_{L^{1}(P_{X})}\right).$
Therefore, all assumptions of [25, Theorem 14] are satisfied under the moment and regularity conditions and we conclude
$W_{1}(P_{X|Y_{\theta_{1}}=y},P_{X|Y_{\theta_{2}}=y})\leq L(y)\|\theta_{1}-\theta_{2}\|,$
i.e., the posterior is locally Lipschitz in the Wasserstein-1 distance with a constant depending on the observation $y$. This verifies Assumption (A3).
@@
We establish a Lipschitz bound on the gradient of the complete-data log-likelihood $\nabla_{\theta}\log p_{X,Y_{\theta}}(y,x)$ with respect to the data variable $x\in\mathbb{R}^{d_{x}}$.
@@
Assumption (A5)
From the discussion of (A3) and (A4) we have
$L(y)\leq C_{1}(1+\|y\|^{2})\quad\text{and}\quad L_{g}(\theta,y)\leq C_{2}(1+\|y\|^{2})$
uniformly over $\theta\in B_{\varepsilon}(\theta^{*})$. Combining these two relations results in
$L(y)\cdot L_{g}(\theta,y)\leq C(1+\|y\|^{2})^{2}.$
We now verify integrability of this upper bound under the marginal distribution $p_{\theta^{*}}$. Note that
$Y=F(X)+\eta,\quad\eta\sim\mathcal{N}\left(0,a^{2}I+b^{2}\mathrm{diag}(F(X)^{2})\right).$
Conditioned on $X=x$, the distribution $Y\mid X=x$ is Gaussian, i.e.,
$Y\mid X=x\sim\mathcal{N}(F(x),a^{2}I+b^{2}\mathrm{diag}(F(x)^{2})).$
Each component $Y_{i}\mid X=x$ has variance at least $a^{2}$ and by the linear growth of $f_{i}$ we get
$\mathrm{Var}(Y_{i}\mid X=x)\leq a^{2}+b^{2}C_{f}^{\prime 2}\|x\|^{2}.$
This implies that $Y\mid X=x$ has sub-Gaussian tails with variance proxy growing at most quadratically in $\|x\|$. Since $X\sim p_{X}$ has finite moments (e.g., Gaussian), the marginal $Y\sim p_{\theta^{*}}$ has finite fourth moment
$\mathbb{E}_{y\sim p_{\theta^{*}}}[\|y\|^{4}]=\int\|y\|^{4}\text{ d}p_{Y_{\theta}^{*}|X=x}(y)\text{ d}p_{X}(x)<\infty.$
Therefore,
$\mathbb{E}_{y\sim p_{\theta^{*}}}[(1+\|y\|^{2})^{2}]<\infty,$
and hence $\gamma(\varepsilon)<\infty$.
@@
With similar arguments and assumptions in (A3) to [27, 25] one can also show local Lipschitz continuity for the total variation distance. Then (A4) is changed, that we need boundedness of the score $\nabla_{\theta}\log p_{X,Y_{\theta}}$ instead of Lipschitz.
@@
While we outlined now conditions under which the convergence is satisfied, it is not clear in practice for which inverse problems they are satisfied. However, preliminary experiments suggest, that the condition $\gamma<c$ can be satisfied in the total variation case. Unfortunately, an analytical study of when this is satisfied is out of reach, due to the non-Gaussianity of the posteriors, as soon as $b>0$. A detailed comparison of the tightness of these analytical bounds of TV and Wasserstein, and the comparison to the true $\gamma$ of the Q-function is left for future work.
@@ 5 Numerical Examples
Algorithm 1 EM Algorithm (Mixed Noise Estimation via Flow Matching)
Input: $\tilde{y}\in\mathbb{R}^{n}$ and initial estimate $\theta_{0}$ by some initial guess $(a,b)$
for r=0,1,,Rr=0,1,...,R do
  E-Step: Simulate $x\sim P_{X}$, and $y\sim P_{Y_{\theta_{r}}|X=x}$
  train the velocity field on the loss $L(w)=\mathbb{E}_{z,x,y,t}[\|v_{t}(y,x_{t})-(x-z)\|^{2}]$
  M-Step: Simulate from the current velocity ODE samples $x\sim P_{X|Y=\tilde{y}}$ for the
  observation $\tilde{y}$
@@
end for

### v2

@@ 3.3 Gradient Descent Based Framework
Under regularity assumptions of (A1) we can swap the following differentiation and integration, and applying subsequently (A2), we get
$\nabla_{1}^{2}\mathcal{Q}(\theta^{\ast},\theta^{\ast})=\mathbb{E}_{y\sim p_{Y_{\theta^{*}}}}\mathbb{E}_{x\sim P_{X|Y_{\theta^{*}}=y}}\left[\nabla_{\theta}^{2}\log p_{X,Y_{\theta^{\ast}}}(x,y)\right]\geq-\mu^{\ast}\mathrm{Id}.$ (7)
@@ 3.4 Step Size Choice Revisited
$\theta^{(k+1)}=\theta^{(k)}+\tau\nabla q(\theta^{(k)}),\qquad\theta^{(0)}\in B_{\epsilon}(\theta^{\ast})$
@@
$0<\sqrt{1-\tau c}+\tau\gamma<1\quad\textrm{for }c\coloneqq\frac{2\mu\lambda}{\lambda+\mu}.$ (8)
@@
For any $\mu,\lambda>0$ and $\gamma$ such that $0\leq\gamma<c=\frac{2\mu\lambda}{\mu+\lambda}$ and $\theta^{0}\in B_{\epsilon}(\theta^{\ast})$ for $\epsilon$ small enough, there exists a step size $\tau\in(0,c^{-1})$ such the gradient ascent iterates in the EM-Algorithm converges to $\theta^{\ast}$.
@@
Since $\mu\geq\lambda$, it holds $c\geq\lambda,$ with equality if and only if $\mu=\lambda$. In particular in the case $\mu>\lambda$, we can relax the constraints on $\gamma$ from [4], i.e. $0\leq\gamma<\lambda$, for convergence to the range $0\leq\gamma<c$ with $c$ from (8). The resulting improvement is illustrated in Figure˜1.
@@ 4 Verification of Assumptions for Mixed Noise Model
  • Regularity and growth of the forward model: Each component $f_{i}:\mathbb{R}^{d}\to\mathbb{R}$ of $F=(f_{1},\ldots,f_{n})^{\top}$ for $i=1,\ldots,n$ is continuously differentiable, bounded, and satisfies the bounded gradient conditions:
    $\|\nabla f_{i}(x)\|\leq C_{f^{\prime}}\quad\text{for all }x\in\mathbb{R}^{d}.$
  • Moment bounds: Existence of first and second moment in $Y$-space under the joint distribution,
    $\int_{\mathbb{R}^{d\times n}}|y|^{k}\,p_{X|Y_{\theta}=y}\,p_{Y_{\theta^{*}}}\text{ d}x\text{ d}y<\infty,$
    for k = 0,1 and 2.
  • Prior regularity: The logarithm of the prior density $p_{X}$ is Lipschitz continuous and has finite moments up to order 4:
@@
the log likelihood is given by
@@
$=-\frac{1}{2}\left(\sum_{i=1}^{n}\log(2\pi\sigma_{i}(x))+\frac{r_{i}(x,y)^{2}}{\sigma_{i}(x)}\right).$
@@
$\log p_{Y_{\theta}|X=x}(y)=-\frac{1}{2}\sum_{i=1}^{n}\left(\log(2\pi\sigma_{i}(x))+\frac{r_{i}(x,y)^{2}}{\sigma_{i}(x)}\right),$
@@
We assume that the required moments exist under the distribution $p_{X|Y_{\theta}=y}\,p_{Y_{\theta^{*}}}$.
Since the log-likelihood is smooth in $\theta\in\Theta$, and the derivatives involve rational expressions in $f_{i}(x)$, $r_{i}(x,y)$, and $\sigma_{i}(x)$, all of which are smooth functions of $x$ and $y$, the gradient and Hessian are continuous. Furthermore, under moment bounds on $y$, the gradient and Hessian norms are integrable, since all involved components are uniformly bounded in $x$ as rational polynomials in $f_{i}(x)$ with nominator of degree larger or equal to the nominators.
For example, let $z=f_{i}(x)\in\mathbb{R}$. Then
$\left|\frac{1}{\sigma_{i}(x)^{2}}-\frac{2r_{i}(x,y)^{2}}{\sigma_{i}(x)^{3}}\right|$ $=\left|\frac{1}{(a^{2}+b^{2}z^{2})^{2}}-\frac{2z^{2}}{(a^{2}+b^{2}z^{2})^{3}}+\frac{4zy_{i}-2y_{i}^{2}}{(a^{2}+b^{2}z^{2})^{3}}\right|$
$\leq\underbrace{\left|\frac{1}{(a^{2}+b^{2}z^{2})^{2}}-\frac{2z^{2}}{(a^{2}+b^{2}z^{2})^{3}}\right|}_{\leq c_{0}<\infty\forall z\in\mathbb{R}}+\underbrace{\left|\frac{4z}{(a^{2}+b^{2}z^{2})^{3}}\right|}_{\leq c_{1}<\infty\forall z\in\mathbb{R}}|y_{i}|+\underbrace{\left|\frac{2}{(a^{2}+b^{2}z^{2})^{3}}\right|}_{\leq c_{2}<\infty\forall z\in\mathbb{R}}|y_{i}|^{2}$
$\leq c_{0}+c_{1}|y_{i}|+c_{2}|y_{i}|^{2}.$
Consequently, for each summand exemplarily for the $\partial_{a^{2}}$ term, we find
$\int_{\mathbb{R}^{d\times n}}\frac{1}{\sigma_{i}(x)^{2}}-\frac{2r_{i}(x,y)^{2}}{\sigma_{i}(x)^{3}}\,p_{X|Y_{\theta}=y}\,p_{Y_{\theta^{*}}}\text{ d}x\text{ d}y\leq\int_{\mathbb{R}^{d\times n}}c_{0}+c_{1}|y_{i}|+c_{2}|y_{i}|^{2}\,p_{X|Y_{\theta}=y}\,p_{Y_{\theta^{*}}}\text{ d}x\text{ d}y<\infty.$
The other bounds for the majorants are obtained analogously, using the gradient condition on the forward operator $f$.
Therefore, the conditions of Assumption (A1)—namely, differentiability of $\theta\mapsto\log p_{X,Y_{\theta}}(x,y)$ and integrable bounds on its gradient and Hessian—are satisfied.
@@
$p_{Y|X=x}(y)=\mathcal{N}(F(x),\bm{\Sigma}(x)),\quad F(x)=(f_{i}(x))_{i=1}^{n},\quad\bm{\Sigma}(x)=\operatorname{diag}(\sigma_{1}(x),\dots,\sigma_{n}(x)).$
@@
$\mathcal{I}(\theta)=\frac{1}{2}\sum_{i=1}^{n}\int_{\mathbb{R}^{d}}\frac{1}{\sigma_{i}(x)^{2}}\begin{pmatrix}1&f_{i}(x)^{2}\\ f_{i}(x)^{2}&f_{i}(x)^{4}\end{pmatrix}p_{X}(x)\,\mathrm{d}x.$
@@
To verify Assumption (A3), we follow the stability result in [30, Theorem 14] in order to obtain Lipschitz continuity of the posterior $\theta\mapsto P_{X|Y_{\theta}=y}$ in the Wasserstein-1 distance.
@@
Since $a^{2},b^{2}$ are bounded, there exists $C_{a^{2}}(y_{i})=C_{a^{2}}(y_{i},a_{\mathrm{min}},a_{\mathrm{max}},b_{\mathrm{min}},b_{\mathrm{max}})>0$ and $C_{b^{2}}(y_{i})=C_{b^{2}}(y_{i},a_{\mathrm{min}},a_{\mathrm{max}},b_{\mathrm{min}},b_{\mathrm{max}})>0$ such that for all $z\in\mathbb{R}$,
$0\leq\left|\frac{1}{(a^{2}+b^{2}z^{2})}-\frac{(y_{i}-z)^{2}}{(a^{2}+b^{2}z^{2})^{2}}\right|<C_{a^{2}}(y_{i})=C_{a^{2}}(y_{i})$
@@
$0\leq\left|z^{2}\left(\frac{1}{(a^{2}+b^{2}z^{2})}-\frac{(y_{i}-z)^{2}}{(a^{2}+b^{2}z^{2})^{2}}\right)\right|<C_{b^{2}}(y_{i})=C_{b^{2}}(y_{i}),$
@@
$\mathrm{maxgrad}\ l(x,y)\leq\sup_{\theta}\frac{1}{2}\sum\limits_{i=1}^{n}\max\{C_{a^{2}}(y_{i}),C_{b^{2}}(y_{i})\}=:C_{\mathrm{grad}}(y).$
@@
for $p=1,2$ constant $L(y)$. The derivation further shows that $L(y)$ depends on $y$ only as a polynomial of degree $2$ of $\|y\|$. Following the proof of [30, Theorem 14], we now obtain that
$W_{1}(P_{X|Y_{\theta_{1}}=y},P_{X|Y_{\theta_{2}}=y})\leq\frac{L(y)\mathbb{E}_{x\sim p_{X}}[\|x\|^{2}]}{\min(Z_{1},Z_{2})}\|\theta_{1}-\theta_{2}\|,$ (9)
where $Z_{1}$ and $Z_{2}$ are the normalization constants $Z_{1},Z_{2}$ of the posterior densities. Note that the assumption $\text{essinf}_{p_{X}}\log p_{Y_{\theta_{1}}|X=x}=0$ in the proof can be avoided by considering $\exp(-[\text{essinf}_{p_{X}}\min(\log p_{Y_{\theta_{1}}|X=x},\\ \log p_{Y_{\theta_{2}}|X=x})]_{-})\leq 1$ in constant.
Finally, the normalization constants can be estimated by
$\min(Z_{1},Z_{2})=\min_{k=1,2}\int\exp(-\log p_{Y_{\theta_{k}}|X=x})p_{X}(x)dx\geq\exp(-C\|y\|^{2})$ (10)
for a constant only depending on $a_{\min},a_{\max},b_{\min},b_{\max}$. Absorbing $\exp(-C\|y\|^{2})$ into $L(y)$ and using the moment assumption on the prior yields that
$W_{1}(P_{X|Y_{\theta_{1}}=y},P_{X|Y_{\theta_{2}}=y})\leq L(y)\|\theta_{1}-\theta_{2}\|$ (11)
for $L(y)$ that is of size $C(\|y\|^{4}+e^{C\|y\|^{2}})$ with a suitable constant $C>0$. This verifies Assumption (A3).
@@
We establish a Lipschitz bound on the gradient of the complete-data log-likelihood $\nabla_{\theta}\log p_{X,Y_{\theta}}(y,x)$ with respect to the data variable $x\in\mathbb{R}^{d}$.
@@
Assumption (A5-SG)
To verify assumption (A5-SG), we first need to show that $Y_{\theta}$ is indeed sub-Gaussian. First note that,
$\mathbb{E}[\|Y_{\theta}\|^{p}]^{1/p}\leq\mathbb{E}[\|F(X)\|^{p}]^{1/p}+\mathbb{E}[\|\eta_{\theta}\|^{p}]^{1/p}.$
Now, due to the boundedness of $F$ and the bound constraints on the parameters $a,b$, there exists a $C\neq C(p)>0$ such that
$\mathbb{E}[\|F(X)\|^{p}]^{1/p}+\mathbb{E}[\|\eta_{\theta}\|^{p}]^{1/p}\leq C\sqrt{p}.$
Then, the characterization of Sub-gaussian properties, in particular the moment property in [35, Proposition 2.6.1], yield the sub-Gaussianity of $Y_{\theta}$. Finally, the bounds in (A3) and (A4) guarantee the existence of the required continuous functions $h_{1}$ and $h_{2}$.
@@
With similar arguments and assumptions in (A3) to [32, 30] one can also show local Lipschitz continuity for the total variation distance. Then (A4) is needs to be exchanged by boundedness of the score $\nabla_{\theta}\log p_{X,Y_{\theta}}$.
@@
Unfortunately, the bounds in the Wasserstein case give an exponential factor, which destroys the integrability required in (A5). Similar to [30, Corollary 19], we only get convergence on sets where $Y$ has bounded radius. This high probability statement does not yield the EM convergence. However, the probabilistic FOS argument gives plausibility to convergence in practice, as will be shown in the next section.
@@ 5 Numerical Examples
Algorithm 1 EM Algorithm (Mixed noise estimation via flow matching, one observation)
1:Input: $\tilde{y}\in\mathbb{R}^{n}$ and initial estimate $\theta_{0}$ by some initial guess $(a,b)$
2:for $r=0,1,\ldots,R$ do
3:  E-Step: Simulate $x\sim P_{X}$, and $y\sim P_{Y_{\theta_{r}}\mid X=x}$
4:  Train the velocity field on the loss
$L(w)=\mathbb{E}_{z,x,y,t}[\|v_{t}(y,x_{t})-(x-z)\|^{2}]$
5:  M-Step: Simulate from the current velocity ODE: $x\sim P_{X\mid Y=\tilde{y}}$
6:  Calculate the loss
@@
7:  Update $\theta$ using gradient descent:
$\theta^{(r+1)}=\theta^{(r)}-\eta\nabla_{\theta}\tilde{L}(\theta)$
8:end for